- Status Report
- August 10, 2022
ISS On-Orbit Status 19 Dec 2002
All ISS systems continue to function nominally, except as noted previously or below.
Wake-up today was at the regular 1:00am EST.
For their first Robotics session today, CDR Ken Bowersox and FE-2/SO Don Pettit were to start out with a review of the DOUG (dynamic operations ubiquitous graphics) software setup. Later, they connected the Lab RWS DCP (robotics workstation display and control panel) power cable, after which the ground powered up the MSS (mobile service system) at 5:00am EST. Start of the 3.5-hr. robotics operations with the SSRMS (space station remote manipulator system) was then scheduled for 9:20am. [The procedure was first to perform a checkout of the Lab RWS DCP auto sequence switch and then maneuver the SSRMS to the W7 MBS PDGF-2 (workstation 7 mobile base system power and data grapple fixture 2) pre-grapple position. From here PDGF-2 was to be grappled, mated, demated, and released twice on the Prime string. After switching to the Redundant string, PDGF-2 was again to be grappled, mated and demated twice, and then released. MCC-H was then to perform normal data dumps to collect SSRMS FMS (force moment sensor) data measured during all of these grapples. Finally, the arm was to be driven to the UHF antenna viewing position to survey the forward end of the antenna.]
FE-1 Nikolai Budarin was tasked to perform a search for a missing Vozdukh cable, probably stowed in the Node.
Budarin also conducted a video survey of IMV (intermodular ventilation) ducts in the Service Module (SM), FGB, DC-1 and Soyuz, with special attention to places where ducts bend, sag, or twist. [The imagery of soft air ducts is part of the search for possible ventilation blockages which would explain the gradual increase of ppCO2 (carbon dioxide partial pressure) noted by ground specialists in the US segment. A procedure for taking airflow measurements is also being developed. This activity will probably be scheduled after the New Year.]
FE-2/SO Pettit activated and checked out EXPRESS Rack 3 (ER3) with the laptop in conjunction with the ground. ER3 was to be powered down again later.
Early in the morning, the crew prepared for and then downlinked a PAO message of greetings for a meeting with Kazakhstan President N. Nazarbaev. The downlink transmission was activated by the automated/preset sequence timer during a Russian comm pass. Subsequently, Budarin recharged the DVCAM (digital video camera).
Later, at 9:00am, Sox, Nikolai and Don were slated for a special television downlink with Christmas messages, to be replayed over the coming holidays. [Wearing Santa caps and scarves, with the little tree they employed during the recent Rockefeller Center downlink message, the crew used scripted texts uplinked last Friday.]
Nikolai Budarin performed lubrication of the blades of the BP (blood pressure) panel compressor used to inflate the BP cuffs of the Russian arterial pressure measuring instrument, a task that was due after one year of BP panel operation.
Sox and Don filled out their second weekly FFQ (food frequency questionnaire), a special software log on the MEC (medical equipment computer) to track nutritional intake.
CDR Bowersox completed the daily routine servicing task of SOSh life support systems maintenance (incl. ASU toilet replacement) and the regular Lab payload status checkup, while FE-1 Budarin performed the daily preparation of the IMS (inventory management system) delta file for downlink to MCC-M/TsUP.
The stripped screw of the ARCTIC-1 refrigerator/freezer cover was successfully removed with the new carbide drill bits on 12/16, by Don Pettit. [The ground now wants to determine whether he was able to torque the remaining seven fasteners to the required 19 in-lbs and stowed the damaged screw in a Ziploc bag in the ARCTIC accessory kit.]
Don was also thanked for doing an excellent job yesterday with the ARIS (active rack isolation system) snubber cap installation with little disturbance of the rack. Too, Ken Bowersox’s EMG (electromyography) calibration dry-run for the Foot (Foot/Ground Reaction Forces During Space Flight) experiment was met with appreciation by the happy Foot Team.
All crewmembers performed their daily physical exercise on TVIS, CEVIS and RED.
The crew is encountering higher cabin temperatures than normal due to the near continuous sunlight in the current high solar Beta angle XPOP attitude. If the cabin becomes too warm, the crew is free to reset the Lab CCAA air conditioner setpoint (thermostat), with notification of the ECLSS (environment control and life support) team on the ground. [The Airlock (A/L) CCAA is being operated without LTL (low temperature loop) flow to minimize the number of wet/dry cycles on the heat exchanger. The heat exchanger coating dissolves into the condensate and is normally transferred into the waste-water tank. However, when there is little to no condensate being collected in the A/L (i.e., no crew in A/L), the water remaining in the water separator evaporates, leaving the minerals behind in the form of scale. This scale caused the old water separator filter to clog, resulting in water being carried over from the heat exchanger and getting into the cabin. ECLSS specialists are working a plan to give the crew relief from the hotter than normal temperatures while maximizing the life on the A/L CCAA hardware.]
The ISS continues to fly in XPOP momentum management attitude mode. MCC-H is prepared to maneuver to +YVV (y-axis in velocity vector) if it is required by Progress battery temperatures.
Begin of sleep time is back at the normal 4:30pm EST.
Today’s targets for the CEO (crew earth observations program) were Inland deltas, S Chad (some of the largest “inland deltas” in the world have developed in southernmost Chad, a major basin on the African continent. Mapping swaths left and right of track were requested of large swamplands [green with brown margins] that are the major visual indicator. The ISS imagery constitutes basic data for a global study of this newly discovered landform type), High Central Andean Glaciers (crew was to photograph near-nadir ice caps, on volcanoes and other ridges. Tropical ice is melting at unprecedented rates and 400/800-mm focal length views hold enough detail for comparative studies), Panama (gap in equatorial cloud. Crew was to look left for an oblique view of the narrowest sector of the landmass), Tuamotu Archipelago (pass over the center of the island chain. Crew to shoot detail of coral reefs. Recent results show that near-vertical handheld views also yield high fidelity subsurface data to a few tens of meters depth), Surface geology, W Libya (oil-rich rock layers are exposed at surface. Views of this remote region from low Earth orbit may reveal structures hard to see from other perspectives. These layers were laid down by glaciers when North Africa straddled the South Pole 450 million years ago), Atlantic storm, Spain (Dynamic event. Major well-formed spiral storm slowly approaching Spain. Wide-angle views will reveal dynamics of interaction of air streams curving into the center of the storm), and Airmass change, NW Atlantic (Dynamic event. Streets of cloud streamers are rising from the sea surface: cold continental air is moistened and warmed from below as it passes out over the NW Atlantic).
CEO images can be viewed at the website
U.S. and Russian Segment Status (as of yesterday).
Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):
- Elektron O2 generator is powered On (24 Amp mode). Vozdukh CO2 scrubber is On. U.S. CDRA CO2 scrubber is Off. TCCS is operational. BMP Harmful Impurities unit: Absorbent bed #1 in Purify mode, bed #2 in Purify mode.
Propulsion System (PS):
- Total propellant load available (SM + FGB + Progress) — 3792 kg (8360 lb) [as of 12/12/02].
Electrical Power Systems (EPS):
- Both P6 channels fully operational. Beta Gimbal Assembly (BGA) 2B and BGA 4B both in dual-angle mode (directed position)..
- SM batteries: Battery #2 in “Cycle” mode. Batteries #5 and 7 disconnected (suspected failed) All other batteries (5) are in “Partial Charge”: mode.
- FGB batteries: Batteries #4 is off line; all other batteries (5) are in “Partial Charge” mode.
- Plasma Contactor Unit PCU-1 is in Standby; PCU-2 is in Standby.
Thermal Control Systems:
- Air conditioner SKV-1 is Off, SKV-2 is Off.
Command & Data Handling Systems:
- C&C-1 MDM is prime, C&C-2 is back-up, and C&C-3 is in standby.
- GNC-1 MDM is prime; GNC-2 is Backup (new patches loaded on both).
- EXT-2 is On (primary), EXT-1 is off.
- LA-1, LA-2 and LA-3 MDMs are all operating.
- PL-1 MDM is On (primary); PL-2 MDM is off (diagnostic
- APS-1 (automated payload switch #1) and APS-2 are both On.
- SM Terminal Computer (TVM): 2 redundant lanes (of 3) operational. Lane 1 is down (as of 11/14).
- SM Central Computer (TsVM): 3 redundant lanes (of 3) operational.
- 3 CMGs on-line (CMG-1 failed).
- State vector — U.S. SIGI-2
- Attitude — U.S. SIGI-2
- Angular rates — U.S. RGA-1
- XPOP 3-CMG Momentum Management (yaw/pitch/roll = 0.52/-8.81/-0.04 deg)
- Solar Beta Angle: -59.4 deg (magnitude increasing)
Communications & Tracking Systems:
- FGB MDM-1 is powered Off; FGB MDM-2 is operating.
- All other Russian communications & tracking systems are nominal.
- S-band is operating nominally.
- Ku-band is operating nominally (except for occasionally parking of antenna due to XPOP increasing thermal temperatures).
- Audio subsystem operating nominally.
- Video subsystem operating nominally.
- MCOR (medium-rate communications outage recorder) is operating nominally.